OpenStudy (anonymous):

Find y'. y=(x^4-3x^2+5)^3

OpenStudy (anonymous):

chain rule: d/dx f(g(x)) = f'(g(x))*g'(x) f(x) = x^3 g(x) = x^4 -3x^2 + 5

OpenStudy (anonymous):

i can walk you through it if you'd like

OpenStudy (anonymous):

would be awesome

OpenStudy (anonymous):

okay so i understand what you have already typed...

OpenStudy (anonymous):

but what do i do with that?

OpenStudy (anonymous):

you'll want to find f'(x) and g'(x). f'(x) = 3x^2 g'(x) = 4x^3 - 6x f'(g(x)) = 3(x^4 -3x^2 + 5)^2 y' = 3(x^4 -3x^2 + 5)^2 * (4x^3 - 6x) = f'(g(x))*g'(x) might seem complicated but it'll come without thinking with practice; i've developed tricks to avoid writing it all out